What is the rate of effusion for a gas that has a molar mass twice that of a gas that effuse at a rate of 3.6 mol/min
1
Expert's answer
2019-11-11T07:49:35-0500

We set the molar masses:

gas A = 1 

gas B = 2

Then assign gas B to r1 (and MM1) gas A to r2 (and MM2) since we want to know the effusion rate for the gas that is twice the molar mass of the other.

Then using Graham's Law:"r1 \/ r2 = (MM2 \/ MM1) ^-1\/2"

r1 / 3.6= (1 / 2) ^-1/2

r1 / 3.6 = 0.70710678

r1 = 2.55 mol/min